Solve the triangle (a is the side opposite α, b is opposite β, and c is opposite γ ). You may estimate to the nearest hundredth but be sure to indicate it is only an estimate, not an exact answer.

1. β=30°, γ=45°, a=5.

2. γ=40°, a=1, b=2

3. a=6, b=2, c=5
Answer #1

@ Given B=30 , 8=450 , a=5 uring the law of shes, xh 450 on From At Brands LAX +LP+ 2y = 180 Lo +30+45º =180 + Set . b = si③ r=40, asl, b=2 using the low of cosiner fa²= 6²+6²-2 be cosA) e²=0²+6²-2 ab cost = 12+1292-841 (2) cos (403) c=1.39 - now